Created by Main.MirkoPojer on 2007-10-17 - 23:15

Sector_5-6_status_161007.jpg:
Comments:
Edit, attach file or add comment
Created by Main.AntonioVergara on 2007-10-17 - 17:02
Sector 8-1
Going-on Activities
- Inner Triplets:
- Left 1: closed, being leak tested
- Right 8: closed, being leak tested
- Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.
Planned activities
- PIC-PC interface test - Wk 43
- Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48)
- ATLAS-RP cable installation - Wks 48-49
- DFBXH:
- Ineo cabling - Wk 43
- Installation of the removed water-cooled cable and RQ6 - Flohe Wk 45
- New cleaning of the DFBXH, the instrumentation will have to be removed and the DFBX re-validated by AT-MEL afterwards - Wk ?
- ElQA
- TP4-A: Inner triplet L1 - as soon as possible since it will be the first time they are performed
- TP4-A: Inner triplet R8 - not before the DFBX is repaired
General
- AUG test on October 19
- Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped.
- A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing
Comments:
Edit, attach file or add comment